OpenStudy (anonymous):

please help ive got an exam tomorrow: determine the equation of the line perpendicular to f(x) = 0.5e^(x+1) at its y-int

jimthompson5910 (jim_thompson5910):

at which point on f(x)?

OpenStudy (anonymous):

at its y-int

jimthompson5910 (jim_thompson5910):

the y-int is when x=0

jimthompson5910 (jim_thompson5910):

plug in x = 0 into f(x) and you get which output value?

OpenStudy (anonymous):

i get 1.36

jimthompson5910 (jim_thompson5910):

you sure? try it again

jimthompson5910 (jim_thompson5910):

oh wait, had a diff value programmed into my calculator for e (weird)

jimthompson5910 (jim_thompson5910):

hold on

OpenStudy (anonymous):

yes i tried it again i still get 1.36

jimthompson5910 (jim_thompson5910):

ok yes I'm now getting that

jimthompson5910 (jim_thompson5910):

so the y-intercept is roughly (0,1.36)

jimthompson5910 (jim_thompson5910):

what is the derivative of f(x) ?

OpenStudy (anonymous):

i get \[f'(x) = (x+1) e ^{x+1}(\frac{ 1 }{ 2 }) \]

OpenStudy (anonymous):

but i'm not sure if its right

jimthompson5910 (jim_thompson5910):

you're using the power rule when you should be using logarithmic differentiation

jimthompson5910 (jim_thompson5910):

have you heard of that method before?

OpenStudy (anonymous):

no...but we have been differentiating log. formulas...i'm guessing that is the method?

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

ok hopefully this page looks familiar to you then (the topics and ideas look familiar) https://www.math.ucdavis.edu/~kouba/CalcOneDIRECTORY/logdiffdirectory/

OpenStudy (anonymous):

so the deriviative would be (x+1)e^x?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

f(x) = 0.5e^(x+1) y = 0.5e^(x+1) ln( y ) = ln( 0.5e^(x+1) ) ln( y ) = (x+1)ln( 0.5e ) ln( y ) = (x+1)( ln( 0.5) + ln( e ) ) ln( y ) = (x+1)( ln( 0.5) + 1 ) d/dx[ ln( y ) ] = d/dx[ (x+1)( ln( 0.5) + 1 ) ] y'/y = ln( 0.5) + 1 y' = y*(ln( 0.5) + 1) y' = 0.5e^(x+1)*(ln( 0.5) + 1) or something along those lines

OpenStudy (whpalmer4):

Huh? since when is \[\frac{d}{dx}[e^{x-1}]\ne e^{x-1}\]?

OpenStudy (whpalmer4):

Sorry, you've got \[\frac{d}{dx}[e^{x+1}]\]but the question stands...

jimthompson5910 (jim_thompson5910):

hmm I guess you could do this f(x) = 0.5e^(x+1) f(x) = 0.5e^x*e^1 f(x) = 0.5e*e^x f ' (x) = 0.5e*e^x

jimthompson5910 (jim_thompson5910):

that's probably a lot easier

OpenStudy (whpalmer4):

\[f'(x) = f(x) = \frac{1}{2}e^{x+1}\]Agreed? Evaluate that at x = 0 to get the instantaneous slope. take the negative reciprocal to find the slope of a perpendicular line, and fit a line through the y-intercept. Am I missing anything?

jimthompson5910 (jim_thompson5910):

no that works too, I just was thinking of something else

jimthompson5910 (jim_thompson5910):

f ' (x) = 0.5e*e^x ---> f ' (x) = 0.5*e^(x+1)

OpenStudy (anonymous):

how is the derivative equal to the same thing as the function?

jimthompson5910 (jim_thompson5910):

but logarithmic differentiation still works here (at least it should)

OpenStudy (whpalmer4):

the magic of e :-)

jimthompson5910 (jim_thompson5910):

derivative of e^x is e^x

OpenStudy (whpalmer4):

it does work of course, if you don't make any errors in the process...

OpenStudy (anonymous):

yes but isn't the derivative of e^bx = (b)e^bx

OpenStudy (anonymous):

or something like that

jimthompson5910 (jim_thompson5910):

you're thinking of the chain rule

jimthompson5910 (jim_thompson5910):

f(x) = e^(g(x)) f ' (x) = g ' (x) *e^(g(x))

jimthompson5910 (jim_thompson5910):

if g(x) = x, then g ' (x) = 1

OpenStudy (anonymous):

so the chain rule isn't applied when e^x+1 correct?

jimthompson5910 (jim_thompson5910):

it can be, but it's not really necessary

OpenStudy (phi):

if you are having a test, you should learn that \[ \frac{d}{dx} e^{f(x)} = e^{f(x)} \frac{d}{dx} f(x) \]

jimthompson5910 (jim_thompson5910):

f(x) = e^(x+1) f ' (x) = e^(x+1) * d/dx[x+1] f ' (x) = e^(x+1) * 1 f ' (x) = 1*e^(x+1) f ' (x) = e^(x+1)

OpenStudy (phi):

in this case you will get back the original e^(x+1) because d/dx (x+1) = 1

OpenStudy (anonymous):

ohhhhhh

OpenStudy (anonymous):

so what about the 0.5...shouldn't the product rule be used?

jimthompson5910 (jim_thompson5910):

0.5 is a constant you can pull out

OpenStudy (phi):

you can put any constant multipliers "out front"

OpenStudy (phi):

like always: d/dx 2x is 2 * d/dx x

OpenStudy (anonymous):

so any number infront of e would be considered a constant?

OpenStudy (phi):

if it is a constant. if it is a variable, you would use the product rule

OpenStudy (anonymous):

ohhh i see. so if there is a cos infron of the e it would be pulled out but if there was a cosx infront of the e than it would be variable and we would use the product rule correct?

OpenStudy (phi):

yes if by cos you mean for example cos(pi) where the angle is constant

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol sorry for all the dumb question but i really need to make this clear for the exam

OpenStudy (phi):

I assume you can finish this problem?

OpenStudy (anonymous):

yes u can thank you both for ur help!

OpenStudy (anonymous):

*i loool awks typo